OpenStudy (anonymous):

Find the indicated derivative. In this case, the independent variable is a (unspecified) function of t. y=(x^100)+(91x^(-1)) find dy/dt

OpenStudy (anonymous):

I have many questions like this, but only need this answered with a detailed example of how you answered it so I can just plug and chug for the other questions! Thanks for your help!

OpenStudy (anonymous):

dy/dt=(the answer)dx/dt

OpenStudy (campbell_st):

well in that case you need the chain rule. as there x = f(t) \[\frac{dy}{dt} = \frac{dy}{dx} \times \frac{dx}{dt}\] so then dy/dt = dy/dx * f'(t)